What is the equation of the perpendicular bisector of the line between the points (2,2) and (6,6)?
Answer by nerdybill(6951)

What is the equation of the perpendicular bisector of the line between the points (2,2) and (6,6)?
.
ref:
http://www.purplemath.com/modules/midpoint.htm
.
The midpoint between the points using the "mid-point formula" is
((x1+x2)/2, (y1+y2)/2)
((2+6)/2, (2+6)/2)
( 8/2, 8/2)
( 4, 4) (this is the mid-point)
.
Now, find the slope of the two points:
m=(y2-y1)/(x2-x1)
m=(6-2)/(6-2)
m=4/4
m=1 (this is the slope of the two points)
.
The perpendicular bisector now has to have a "negative reciprocal" of 1 so our NEW line should have a slope of -1.
.
So, our new line has:
slope = -1
crossing (4,4)
.
y=mx + b
plugging in our info:
4=(-1)4 + b
4= -4 + b
8 = b
.
Finally, we have:
m = -1
b = 8
so,
y = -x + 8 (THIS is what they're looking for)